Sole the equation x^2=3-2x by completing the square

Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)



"Sole" the equation? Is that "sole" as in the fish or as in the bottom part of a shoe?

If you actually want to "solve" the equation by completing the square, read on.

1. Put the terms with variables in the LHS, and leave the constant term in the RHS.



2. Divide the coefficient on the first degree term by 2, square the result, and add that result to both sides of the equation. 2 divided by 2 is 1, 1 squared is 1, add 1 to both sides.



3. Factor the perfect square polynomial in the LHS:



4. Take the square root of both sides (Remember both positive and negative roots)



5. Solve both linear equations:





John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by stanbon(75887)

Solve the equation x^2=3-2x by completing the square
----------
x^2 + 2x + ? = 3 + ?
---------
x^2 + 2x + 1 = 3 + 1
---------
(x+1)^2 = 4
=========
x + 1 = 2 or x+1 = -2
x = 1 or x = -3
===================
Cheers,
Stan H.
==========================